OpenStudy (anonymous):

12 persons are seated around a round table. What is the probability that two particular persons sit together ?

OpenStudy (astrophysics):

Lets draw a diagram first  looks like the sun! Anyways, so probability is just \[Probability = \frac{ \text{number of favourable outcomes} }{\text {number of possible outcomes} }\]

OpenStudy (anonymous):

possible outcomes is 11!

OpenStudy (anonymous):

No. of favorable outcomes?

OpenStudy (astrophysics):

It says, that two persons sit together, 11! is the number of ways they can sit around the table but it's not for two people right, so we have 10! instead

OpenStudy (astrophysics):

And number of possible outcomes is 11

OpenStudy (anonymous):

But answer is 2/11

OpenStudy (astrophysics):

11!*

OpenStudy (astrophysics):

Yes so \[\frac{ 10! \times 2! }{ 11! }\]

OpenStudy (anonymous):

Oh okay. thank you @Astrophysics :)

OpenStudy (anonymous):